### v1

# The Hidden Link Between RLHF and Contrastive Learning

Abstract
Alignment of large language models (LLMs) with human values has recently garnered significant attention, with prominent examples including the canonical yet costly Reinforcement Learning from Human Feedback (RLHF) and the simple Direct Preference Optimization (DPO). In this work, we demonstrate that both RLHF and DPO can be interpreted from the perspective of mutual information (MI) maximization, uncovering a profound connection to contrastive learning. Within this framework, both RLHF and DPO can be interpreted as methods that performing contrastive learning based on the positive and negative samples derived from base model, leveraging the Donsker–Varadhan (DV) lower bound on MI (equivalently, the MINE estimator). Such paradigm further illuminates why RLHF may not intrinsically incentivize reasoning capacities in LLMs beyond what is already present in the base model. Building on the perspective, we replace the DV/MINE bound with the Jensen–Shannon (JS) MI estimator and propose the Mutual Information Optimization (MIO). Comprehensive theoretical analysis and extensive empirical evaluations demonstrate that MIO mitigates the late-stage decline in chosen-likelihood observed in DPO, achieving competitive or superior performance across various challenging reasoning and mathematical benchmarks. The code is available at: https://anonymous.4open.science/r/MIO-63E6/
1 Introduction
Recently, large language models (LLMs) have attained state-of-the-art performance on a wide array of natural-language tasks [1]. Aligning their outputs with human value is commonly achieved through the methodology of reinforcement learning from human feedback (RLHF) [2][3], whose fine-tuning recipe has become one of the industry standards. However, its dependence on two step reinforcement learning presents challenges [4], such as computational inefficiency and training instability. To mitigate these limitations, alternative one-step approaches such as direct preference optimization (DPO) [5] and its variants [6][7] have been proposed , which has successfully eliminated the need for explicit reward modeling. Specifically, an implicit reward based on the likelihood of preference data is defined, which results in significant gains in efficiency while preserving competitive performance.
Major Questions 1. Could RLHF be theoretically framed as a form of contrastive learning? 2. Does RLHF fundamentally extend a base model’s capabilities or only amplify existing ones? 3. How can we overcome DPO’s limitations to further improve the reasoning and mathematical performance of fine-tuned LLMs?
Several recent works have hinted at conceptual links between RLHF and imitation learning [8]. DIL [4] show that optimizing a policy with RLHF implicitly maximizes the likelihood of human-preferred outputs, formally proves that RLHF for aligning language models could be view as an imitation learning problem. IRL[9] examines the pretraining and fine-tuning of LLMs from an imitation learning standpoint, argue that standard next-token fine-tuining is a form of imitation learning. Beyond imitation learning, several works have drawn intuitive parallels between Direct Preference Optimization (DPO) and contrastive learning [10]. For example, Xu et al. [11] and Chen et al. [12] treat DPO as a form of contrastive learning; however, their analysis still lacks a formal theoretical foundation. Such observation motivates our central research question: Is there a unified theoretical framework that rigorously establishes the connection between RLHF and contrastive learning?
Figure 1: RLHF performs as contrastive learning: PPO/DPO pushes the base model toward the positive model $\pi_{\text{chosen}}$ and away from the negative model $\pi_{\text{rejection}}$. Search trees are generated by repeatedly sampling from the models for a given problem. The gray indicates paths that are unlikely to be sampled by the model, while the black indicates paths that are likely to be sampled. “+1” indicates the correct paths, which have positive rewards. The base model contains both the correct and false path. The positive model is more likely to reach the correct path, whereas the negative model is more likely to follow incorrect paths. After the RLHF process, the policy model becomes closer to the positive model and diverges from the negative model.
Recently, it is widely believed that Reinforcement Learning with Variable Rewards (RLVR), a variant of RLHF, facilitates the continual self-improvement of LLMs, thereby enabling them to acquire novel reasoning capabilities beyond those of their corresponding base models [13]. However, recent empirical findings challenge this assumption. Specifically, Yue et al. [14] reports that RLVR training does not, in fact, induce fundamentally new reasoning trajectories. Their analysis reveals that the reasoning paths exhibited by RL-trained models largely fall within the sampling distribution of the base models, implying that the reasoning capabilities demonstrated post-RL training may already be latent in the base models prior to fine-tuning. Nevertheless, the observed phenomenon in this work lacks a rigorous theoretical explanation. This brings us to our second research question: Is there a theoretical framework that can formally establish whether RLHF truly extends the capabilities of a base model, or merely amplifies its pre-existing capacities?
RLHF and its variants fundamentally adhere to a reward maximization objective, often instantiated through parametric models such as the Bradley–Terry (BT) model [15]. This modeling paradigm, while effective in aligning with preference data, has been shown to suffer from overfitting [16, 17], frequently leading to suboptimal generalization and misalignment with true user preferences [18, 19]. Recent empirical studies further indicate that DPO and its variants progressively focus on unlearning the rejected responses. This process inadvertently increases the model’s tendency to generate out-of-distribution responses, rather than reinforcing alignment with the chosen responses [18]. From a theoretical perspective, Yan et al. [20] analyzes the optimization dynamics of DPO and demonstrates that the shared tokens between chosen and rejected responses lead to gradient interference. This interference introduces instability and a coupled degradation, whereby suppressing rejected responses inadvertently reduces the model’s confidence in the selected responses. Such simultaneous unlearning of both rejected and chosen outputs ultimately degrades downstream performance—particularly in reasoning-intensive and mathematical tasks [17, 16, 6]. This observation motivates our third research question: Can we design a principled algorithm that overcomes inherent shortcomings of DPO, while effectively enhancing model performance on complex reasoning and mathematical tasks?
This paper makes the following contributions: (1) We present a unified mutual-information perspective on RLHF and contrastive learning, demonstrating that both RLHF and DPO optimize a Donsker–Varadhan (DV) [21, 22] lower bound on mutual information, thereby aligning with contrastive objectives. (2) This framework provides a theoretical explanation for why RLHF fails to elicit novel reasoning trajectories: it can only amplify reasoning patterns already present in the base model. (3) To overcome the degradation issue in DPO, we propose the Mutual Information Optimisation (MIO), a one-step algorithm based on a Jensen–Shannon MI estimator [23], which achieves state-of-the-art performance on multiple reasoning and mathematical benchmarks. (4) We further provide a rigorous theoretical analysis showing that MIO preserves meaningful learning signals on chosen responses, even as rejected responses are suppressed.
2 Rethinking RLHF: A Contrastive Learning Perspective
Answer to Question 1: RLHF is a form of Contrastive Learning in Mutual Information frame.
Given a prompt $x=[\text{x}_{1},\text{x}_{2},...]$, we are presented with two candidate responses: $y_{w}=[\text{y}_{1},\text{y}_{2},...]$, the response preferred by humans; and $y_{l}$, the response with lower preference. Our goal is to fine-tuning the language model, enhancing its propensity to generate responses that are consistent with human preferences. Firstly, we assume that all positive examples $y_{w}$ are drawn from a preferred language model $\pi_{\text{chosen}}$, while negative examples $y_{l}$ are drawn from a dispreferred language model $\pi_{\text{rejection}}$. We assume that both $\pi_{\text{chosen}}$ and $\pi_{\text{rejection}}$ are energy-based models (EBMs) [24], as discussed in Appendix A. Our objective is to adjust the target model $\pi_{\theta}$ to maximize mutual information with ${\pi}_{\text{chosen}}$ while minimizing mutual information with $\pi_{\text{rejection}}$.
We further reformulate this problem as a contrastive learning objective:
$\max_{\theta}\,I(\pi_{\theta},\pi_{\text{chosen}})\qquad\text{and}\qquad\min_{\theta}\,I(\pi_{\theta},\pi_{\text{rejection}})$ (1)
We begin by considering the unconstrained maximization of mutual information with respect to $\pi_{\text{chosen}}$, which reduces to a standard imitation learning problem:
$\max_{\theta}\,I(\pi_{\text{chosen}},\pi_{\theta})$ (2)
In Appendix B, we show that the contrastive learning objective and the imitation learning objective are equivalent, since increasing one objective decreases the other.
Since mutual information (MI) cannot be directly computed, we introduce an estimator for the lower bound of mutual information, namely the Mutual Information Neural Estimation (MINE) [22]:
$I(\pi_{\theta},\pi_{\text{chosen}})=\sup_{\phi}{}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}[T_{\phi}]-\log\mathbb{E}_{P_{\pi_{\theta}}P_{\pi_{\text{chosen}}}}[e^{T_{\phi}}],$ (3)
where $P_{\pi_{\theta}\pi_{\text{chosen}}}$ denotes the joint distribution of $\pi_{\theta}$ and $\pi_{\text{chosen}}$, while $P_{\pi_{\theta}}P_{\pi_{\text{chosen}}}$ represents the marginal distribution of them. MINE maximizes the mutual information between $\pi_{\theta}$ and $\pi_{\text{chosen}}$ by optimizing a network-based function $T_{\phi}$.
Hence, we can derive the following expression for mutual information (see Appendix C.1):
$I(\pi_{\theta},\pi_{\text{chosen}})=\sup_{\phi}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}[T_{\phi}]-\log\left(\mathbb{E}_{P_{\pi_{\theta}}}\left[\mathbb{E}_{P_{\pi_{\text{chosen}}}}[e^{T_{\phi}}]\right]\right)$ (4)
To improve the optimization, we introduce a lower bound for the joint and marginal distributions:
$I(\pi_{\theta},\pi_{\text{chosen}})\geq\sup_{\phi}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}[T_{\phi}]-\log\left(\mathbb{E}_{P_{\pi_{\theta}}}\left[2\mathbb{E}_{P_{\bar{\pi}}}[e^{T_{\phi}}]\right]\right),$ (5)
where $\bar{\pi}(\cdot|x)=\frac{1}{2}\pi_{\text{chosen}}(\cdot|x)+\frac{1}{2}\pi_{\text{rejection}}(\cdot|x)$. Considering the following Monte Carlo estimator:
$\mathbb{E}_{\pi_{\theta},\bar{\pi}}\left[e^{T_{\phi}(\pi_{\theta},\bar{\pi})}\right]\approx\frac{1}{2}\left(\frac{1}{M}\sum_{i=1}^{M}e^{T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})}+\frac{1}{N}\sum_{i=1}^{N}e^{T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})}\right).$ (6)
We substitute Equation (6) into Equation (5), can estimate the lower bound of mutual information as:
$\max_{\theta}\hat{I}=\log\left(\frac{e^{T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})}}{\frac{1}{M}\sum_{i=1}^{M}e^{T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})}+\frac{1}{N}\sum_{i=1}^{N}e^{T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})}}\right).$ (7)
When $M=N=1$, this objective could be simplified to:
$\max_{\theta}\log\sigma\left(T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})-T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})\right),$ (8)
where $T_{\phi}$ serves as a parameterized model used to estimate mutual information between two models, and is equivalent to the reward model in RLHF, see in Appendix A. Then, by treating $T_{\phi}$ as the reward model, we can derive an objective function that aligns precisely with the reward loss under the BT preference assumption in RLHF.
Furthermore, if we select $T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})=\log\frac{\pi_{\theta}}{\pi_{\text{chosen}}}$, and substitute it into Equation (8), we can arrive at the objective function of Direct Preference Optimization (DPO) [5]:
$\max_{\theta}\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)$ (9)
Keep $T_{\phi}(\pi_{\theta},\pi)=\log\frac{\pi_{\theta}}{\pi}$. Only substituting this into the right-hand side of the MINE objective (4) and applying certain approximations yields the following objective used in the second stage of RLHF, where the policy model is updated using the reward model[19, 25]:
$\max\mathbb{E}_{Joint}[T_{\phi}]-\mathrm{KL}(\pi_{\theta}\,\|\,\pi_{\text{ref}}).$ (10)
The approximation error introduced here is upper-bounded (see Appendix D), and the bound depends on the ratio $\frac{\pi_{\theta}(\cdot\mid x)}{\pi_{\text{ref}}(\cdot\mid x)}$: the smaller this ratio, the smaller the approximation error. This also explains why PPO introduces a clipping function to constrain the update magnitude of the policy—large deviations in this ratio lead to large approximation errors, which in practice degrade performance. As the derivatives of $\hat{I}(\pi_{\theta},\pi_{\text{chosen}})$ and $\hat{I}(\pi_{\theta},\pi_{\text{rejection}})$ are opposites, maximizing $\hat{I}(\pi_{\text{chosen}},\pi_{\theta})$ simultaneously minimizes $\hat{I}(\pi_{\text{rejection}},\pi_{\theta})$. This shows that optimizing the mutual information between $\pi_{\theta}$ and $\pi_{\text{chosen}}$ inherently minimizes the mutual information between $\pi_{\theta}$ and $\pi_{\text{rejection}}$, thereby transforming the task into a contrastive learning framework.
Thus, we have derived the DPO objective function from a contrastive learning perspective by introducing a mutual information neural estimator (MINE), illustrating that DPO is essentially a special case of contrastive learning.
Furthermore, through the selection of $T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})=\log\frac{\pi_{\theta}}{\pi_{\text{chosen}}}$, we can draw some insights:
  1. 1.
    Reward learning in RLHF is equivalent to the task of searching for a function $T\in\mathcal{F}_{\text{all}}$ within a specified function space. This function $T$ minimizes the discrepancy between estimated and actual unobservable mutual information (see Appendix C), thereby making the estimated and actual mutual information as close as possible.
  2. 2.
    After learning the reward model in RLHF, the subsequent update of the policy model is equivalent to a contrastive learning task. This phase can be seen as maximizing the mutual information between $\pi_{\theta}$ and the positive sample $\pi_{\text{chosen}}$, while minimizing the mutual information with the negative sample $\pi_{\text{rejection}}$. If we limit the search for $T$ to a subset $\mathcal{F}_{\text{sub}}$, (see Appendix C.4), we can obtain the DPO loss function.
Answer to Question 2: RLHF does not fundamentally extend a base model’s capabilities.
Definition 2.1 (Model Capacity).
Given an input question $x$ and its correct answer $y_{w}$, we say that a model $\pi_{\theta}$ has the capacity to correctly answer the question if $\pi_{\theta}(y_{w}\mid x)\neq 0$; that is, the model assigns non-zero probability to the correct answer.
Corollary 2.1.
RLHF does not inherently incentivize the emergence of reasoning capabilities in large language models that go beyond those already present in the base model.
In practice, the base model is typically frozen and used as a reference model. In this setting, positive samples are drawn from $\pi_{\text{chosen}}$, which can be interpreted as an enhanced version of the base model, while negative samples represent a degraded variant. Consequently, the optimal policy $\pi_{\theta}$ is expected to fully align with the $\pi_{\text{chosen}}$. However, since $\pi_{\text{chosen}}$ is defined as an energy-based reweighting of $\pi_{\text{reference}}$, its support is inherently limited by that of the base model. If the base model assigns zero probability to a correct reasoning path (i.e., $\pi_{\text{reference}}(y_{w}\mid x)=0$), then even a well-trained reward model cannot recover or induce such reasoning capability in $\pi_{\theta}$.
RLHF does not fundamentally extend a base model’s capabilities; rather, it amplifies the reasoning abilities already present within base model. It refines and reinforces these existing patterns but does not create entirely new capabilities. This provides a theoretical explanation for experimental findings in some studies [14], which observe that RLHF does not enhance the capabilities of base model.
3 Beyond DPO: From DV to JSD Mutual-Information Optimisation
3.1 Why should JSD be used instead of MINE in Alignment?
Subsequent research has shown that the DV-based objective used in MINE suffers from significant instability in practice [26]. Specifically, its variance can grow exponentially with the true mutual information (MI) [27], leading to exploding gradients and requiring very large batch sizes to maintain stable training [28]. Poole et al. [26] further demonstrate that DV-style bounds like MINE exhibit low bias but extremely high variance, causing erratic gradient updates.
In the work [29], it is emphasized that accurate MI estimation is intrinsically challenging: high bias or high variance can lead to overfitting of spurious correlations. Thus, in contrastive learning, stable MI lower bounds (e.g., MINE or the Jensen–Shannon (JS) bound) should be prioritized over maximizing numerical MI estimates.
Additionally, Sordoni et al. [30] point out that the MINE bound is upper-bounded by $\log K$, where $K$ is the number of negative samples. When the true MI greatly exceeds $\log K$, the bound severely underestimates MI.
Collectively, these studies [28, 29] highlight that in regimes with few negative samples, MINE suffers from prohibitively large variance [26], leading to unstable gradients. Therefore, replacing MINE with a variance-reduced alternative, such as the JSD estimator [31], is crucial for reliable optimization.
As our primary objective is to maximize MI rather than precisely estimate its value [23], we can rely on non-KL divergences, which offer favorable trade-offs. For instance, the Jensen-Shannon MI estimator [32]—similar to the binary cross-entropy used in minimizing total correlation [33]—is well-understood in neural network optimization and is more stable in practice compared to the DV-based objective [23]:
$I_{\text{JSD}}(\pi_{\theta};\pi_{\text{chosen}}):=\sup_{\theta\in\Theta}\mathbb{E}_{\pi_{\theta}}\left[-\text{sp}(-T_{\phi}(\pi_{\theta},\pi_{\text{chosen}}))-\mathbb{E}_{\bar{\pi}}[\text{sp}(T_{\phi}(\pi_{\theta},\bar{\pi}))\right].$ (11)
We investigate the behavior of mutual information (MI) estimators in the low-negative-sample regime using a synthetic bivariate Gaussian setting, where the correlation coefficient $\rho$ controls the dependency between variables. The ground-truth MI is given analytically as $I(X,Y)=-\frac{1}{2}\log(1-{\rho}^{2})$.
Figure 2: Left: Estimated mutual information under the bivariate Gaussian setting with varying $\rho$, using MINE and JSD estimators with $K=1$ negative sample. While both methods track the ground-truth trend, JSD is consistently than MINE. Right: Gradient variance of each estimator. The JSD estimator yields substantially lower variance, indicating more stable optimization behavior under limited negative sampling.
The left panel of Figure 2 compares the MINE and JSD estimators with a single negative sample ($K=1$). While both estimators qualitatively capture the ground-truth MI trend, JSD does not consistently outperform MINE in terms of accuracy. However, when gradient stability is prioritized over precise MI estimation, JSD demonstrates a clear advantage [34]. To further investigate this, we analyze the gradient variance induced by each estimator, with the right panel of Figure 2 showing the gradient variance computed from the first-layer weights of a three-layer MLP trained for each estimator’s objective. Under conditions with limited negative samples, JSD exhibits significantly greater stability than MINE, making it particularly suitable for training objectives, such as Alignment, that rely on contrastive MI estimates with sparse negative sampling [12].
3.2 From DPO to Mutual-Information Optimisation (MIO)
Considering the following Monte Carlo estimator:
$\begin{split}\mathbb{E}_{\bar{\pi}}\big{[}\text{sp}(T_{\phi}(\pi_{\theta},\bar{\pi})\big{]}\approx\frac{1}{2}\Big{(}&\frac{1}{M}\sum_{i=1}^{M}\log\big{(}1+\exp(T_{\phi}(\pi_{\theta},\pi_{\text{chosen}}))\big{)}\\ +&\frac{1}{N}\sum_{i=1}^{N}\log\big{(}1+\exp(T_{\phi}(\pi_{\theta},\pi_{\text{rejection}}))\big{)}\Big{)}.\end{split}$ (12)
By replacing the DV/MINE [22] surrogate in Section 2 with the JS-based surrogate given in Equations (11) and (4), and substituting  (12) into  (11), we obtain the following empirical objective:
$\begin{split}\hat{I}_{\mathrm{JSD}}=-&\frac{1}{M}\sum_{i=1}^{M}\!\log\Big{(}1+e^{-T_{\phi}(\pi_{\theta},\pi_{\text{c}})}\Big{)}\\ -\frac{1}{2}\Biggl{[}&\frac{1}{M}\sum_{i=1}^{M}\!\log\Bigl{(}1+e^{\;T_{\phi}(\pi_{\theta},\pi_{\text{c}})}\Bigr{)}+\frac{1}{N}\sum_{j=1}^{N}\!\log\Bigl{(}1+e^{\;T_{\phi}(\pi_{\theta},\pi_{\text{r}})}\Bigr{)}\Biggr{]}.\end{split}$ (13)
Setting $M=N=1$ and constraining the critic to $T_{\phi}(\pi,\pi_{\theta})=\log\frac{\pi_{\theta}}{\pi}$ gives a closed-form loss, which we call Mutual-Information Optimisation (MIO):
$\mathcal{L}_{\text{MIO}}(\theta)\;=\;\log\!\Bigl{(}1+\tfrac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\theta}(y_{w}\mid x)}\Bigr{)}+\frac{1}{2}\log\!\Bigl{(}1+\tfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}\Bigr{)}+\frac{1}{2}\log\!\Bigl{(}1+\tfrac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\Bigr{)}.$ (14)
Why MIO is Preferable to DPO Under Low $K$.
With $K=1$ negative sample, the DV/MINE surrogate induces unstable gradients due to its high-variance denominator, whereas the JSD surrogate reduces gradient variance without compromising performance (Fig. 2, right panel). Since DPO is based on DV/MINE [22] and MIO is based on JSD [23], the policy update driven by (14) is more stable, particularly when negative comparisons per prompt are limited [12]. When few negative samples are available, the gradients from MINE become unstable, with the choice of negatives significantly affecting the estimate and amplifying gradient variance. Given that our primary goal is to obtain reliable gradients rather than precise MI estimates, the JSD estimator is preferred [34]. Due to its substantially lower variance, MIO results in smoother optimization trajectories and mitigates gradient oscillations caused by noisy negative sampling.Empirical results in Figure 5 confirm that MIO strengthens alignment with the chosen responses while reducing consistency with the rejected responses. In contrast, DPO [35] decreases consistency with both the chosen and rejected responses, which may impair the model’s reasoning and mathematical performance [17, 16].
4 Experiment
4.1 DPO Breaks When Rejection near 0, But MIO Not!
Following [20], the toy model setup, shown in Figure 3, consists of a discrete space with 4 prompts and 10 responses. The policy $\pi_{\theta}$, implemented as a three-layer MLP, processes a one-hot input vector and outputs a categorical distribution over the responses, which are divided into three categories: selected responses (first 4 dimensions), rejected responses (next 4 dimensions), and unseen responses (final 2 dimensions). Each prompt has a corresponding optimal response (e.g., response 1 is optimal for prompt 1). In DPO or MIO training, preference data is generated through a mini-batch sampling strategy, where an ideal annotator matches each prompt with its optimal response, creating a diagonal preference matrix. In each mini-batch, one additional response is randomly selected to form preference pairs, ensuring diverse gradient updates.
Figure 3: Toy model setup[20]. Left: the optimal policy where the highlighted blocks represent optimal responses. Right: preference dataset construction.
Figure 4: Overview of the DPO and MIO dynamics.From left to right, the figures show the initial state and the likelihood dynamics for chosen/rejected/unseen responses in Scenarios 1 to 4, similar to the left diagram in Figure 2: (1) both chosen and rejected responses are very small, (2) chosen is normal and rejected is very small, (3) chosen is small and rejected is normal, and (4) both chosen and rejected is normal.
Figure 4 contrasts the learning dynamics of DPO with those of our proposed MIO objective. When the probabilities of both the chosen and rejected responses remain in a moderate range, DPO converges smoothly and the likelihood of the chosen response declines only marginally relative to that of the rejected one. Once the rejected–response probability becomes very small, however, DPO develops a pathological synchronous collapse: the probabilities of both the chosen and the rejected responses decrease in lock-step. Across all regimes we tested, MIO displays no such joint degradation. When the rejected probability is pushed towards zero, DPO and its derivatives exhibit synchronous decline of the chosen and rejected likelihoods, mirroring what is observed in practice [20, 12, 8], whereas MIO remains entirely immune to this failure mode. A formal explanation of this qualitative difference is provided in Section 5.
The toy model serves as an abstract simulation that amplifies the 3D-properties of DPO [20], which leads to a decrease in the likelihood of both chosen and rejected responses. However, MIO does not exhibit this effect. While the toy model setup differs significantly from real-world LLM training—such as in sampling frequency—it provides valuable insights. In real-world scenarios, DPO and MIO are typically trained over one epoch, with each data point used only a few times. In contrast, the toy model samples the same data points repeatedly. Conceptually, this is akin to treating each input/output as a token rather than a full prompt/response, where each token may be sampled multiple times during real-world training. The code will be provided in the supplementary materials.
4.2 Experimental Setup
We fine-tune the Mistral-7B-Base [36] on the 64 k-prompt UltraFeedback-Binarized [37] corpus, where each prompt is paired with a chosen and a rejected completion selected by GPT-4, yielding 64 k positive–negative pairs; we retain the official train/validation split of UltraFeedback-Binarized.
Model quality is assessed on eight public mathematical and reasoning suites: Hendrycks MATH [38] (12.5 k competition-level problems plus the level-5 “hard” subset), Minerva Math [39] (quantitative-reasoning items distilled from GSM8K and MATH), MultiMedQA [40] (six medical QA benchmarks spanning USMLE, PubMedQA, and consumer health queries), MathQA [41] (37 k multiple-choice word problems with executable rationales), GSM8K [42] (8.5 k grade-school arithmetic questions), AQuA-RAT [43] (100 k algebraic problems with free-text solutions), and MuSR [44] (narrative tasks requiring multistep soft reasoning over 700–1000-word stories). The code, hyperparameter settings, and evaluation metrics will be provided in the supplementary materials.
Table 1: Evaluation results on mathematical & reasoning benchmarks (the higher, the better). The best and second best numbers are marked with boldface and underline.
Method ($\downarrow$) / Benchmark ($\rightarrow$) Hendrycks Math Minerva Math Multimedqa MathQA GSM8K Aqua Rat Math Hard MUSR
SFT 0.14 8.80 49.03 29.91 32.29 19.29 2.72 39.94
DIL[4] 0.04 6.46 49.54 28.94 28.35 21.33 2.11 38.36
DPO[5] 0.04 8.14 49.21 30.72 36.47 19.69 2.72 43.65
IPO[45] 0.08 6.50 48.39 29.41 3.71 19.69 0.23 35.71
KTO[46] 0.04 10.38 49.76 30.56 41.75 21.65 3.43 43.39
NCA[12] 0.14 8.92 49.58 30.59 37.68 18.50 2.49 39.68
ORPO[47] 0.04 9.68 49.40 29.78 36.39 24.80 2.57 42.06
SimPO[6] 0.19 8.36 49.16 30.93 39.20 18.11 2.49 39.02
SLiC[48] 0.08 8.72 49.74 30.66 39.73 20.47 3.17 42.99
MIO 0.21 9.08 49.82 30.96 40.16 24.83 3.57 43.42
Across the eight mathematical and reasoning benchmarks, MIO achieves the best performance on five tasks and ranks second on two others. These results demonstrate that MIO offers superior generalization in reasoning and mathematical problem-solving compared to existing alignment methods.
4.3 MIO prevents the chosen-reward collapse
Figure 5: Training curves on Mistral-7B-Base. Unlike DPO, IPO, and ORPO, MIO increases the rewards of chosen responses while suppressing rejected ones, thereby avoiding the “synchronous collapse” observed in prior methods.
During DPO training, the rewards of all responses ($y_{w}$ and $y_{l}$) often decrease simultaneously, a harmful “unlearning” effect that degrades downstream performance [4, 6, 17]. Figure 5 shows that the proposed MIO objective completely eliminates this failure mode: the chosen-rewards increase while the rejected rewards drop slightly. Section 5 provides a theoretical explanation of this behavior.
5 Why MIO Outperforms DPO?
5.1 Failure mode of DPO
For a preference triple $(x,a^{+},a^{-})$ the DPO loss can be written as [20]:
$\ell_{\mathrm{DPO}}=\log\!\Bigl{(}1+\bigl{[}\alpha\,z^{\beta}\bigr{]}\Bigr{)},\qquad\alpha=\Bigl{[}\tfrac{\pi_{0}(a^{+}\!\mid x)}{\pi_{0}(a^{-}\!\mid x)}\Bigr{]}^{\beta},\;z=\tfrac{\pi_{\theta}(a^{-}\!\mid x)}{\pi_{\theta}(a^{+}\!\mid x)}.$
Denoting $\pi^{+}\!=\!\pi_{\theta}(a^{+}\!\mid x)$ and $\pi^{-}\!=\!\pi_{\theta}(a^{-}\!\mid x)$, differentiation gives
$\frac{\partial\ell_{\mathrm{DPO}}}{\partial\pi^{+}}=-\frac{\alpha\beta}{1+\alpha z^{\beta}}\frac{z^{\beta}}{\pi^{+}},\qquad\frac{\partial\ell_{\mathrm{DPO}}}{\partial\pi^{-}}=\frac{\alpha\beta}{1+\alpha z^{\beta}}\frac{z^{\beta-1}}{\pi^{+}}.$
Hence, we have that $\lvert\partial_{\pi^{-}}\ell\rvert/\lvert\partial_{\pi^{+}}\ell\rvert=\pi^{+}/\pi^{-}$. The gradient on the rejected probability therefore grows much faster than that on the chosen probability; as $\pi^{-}\!\to 0$ we have $\partial_{\pi^{-}}\ell\!\to\!\infty$ but $\partial_{\pi^{+}}\ell\!\to 0$. Because chosen and rejected completions usually share many tokens, the vanishing positive gradient cannot offset the dominant negative gradient, so the model is driven to lower both probabilities. This synchronous likelihood collapse—observed empirically in late-stage DPO training—degrades overall performance [17].
5.2 Stability of MIO
Answer to Question 3: MIO leads to a more stable training process!
The MIO objective for the same triple is
$\ell_{\mathrm{MIO}}=\ln\!\bigl{(}1+e^{-\beta\operatorname{LR}^{+}}\bigr{)}+\tfrac{1}{2}\ln\!\bigl{(}1+e^{\beta\operatorname{LR}^{+}}\bigr{)}+\tfrac{1}{2}\ln\!\bigl{(}1+e^{\beta\operatorname{LR}^{-}}\bigr{)},$
with $\operatorname{LR}^{\pm}=\log\pi^{\pm}-\log\pi_{0}^{\pm}$. Writing $\sigma^{\pm}=\sigma\!\bigl{(}\beta\operatorname{LR}^{\pm}\bigr{)}$, one obtains
$\partial_{\pi^{+}}\ell_{\mathrm{MIO}}=\frac{\beta}{\pi^{+}}\!\bigl{(}1.5\sigma^{+}-1\bigr{)},\qquad\partial_{\pi^{-}}\ell_{\mathrm{MIO}}=\frac{\beta}{2\pi^{-}}\sigma^{-}.$
Proposition 5.1 (Selective suppression of negatives).
Let $\pi^{-}\!\to 0$. Then, we have:
$\left\lvert\frac{\partial\ell_{\mathrm{MIO}}}{\partial\pi^{-}}\right\rvert=\frac{\beta}{2\pi^{-}}\sigma^{-}\;\longrightarrow\;\infty,\qquad\left\lvert\frac{\partial\ell_{\mathrm{MIO}}}{\partial\pi^{+}}\right\rvert=\frac{\beta}{\pi^{+}}\bigl{\lvert}1.5\sigma^{+}-1\bigr{\rvert},$
so the negative likelihood continues to be forcefully suppressed, while the positive gradient remains bounded and therefore does not vanish. Consequently, MIO preserves a strong corrective signal for rejected responses without impairing the optimisation of chosen responses.
Proposition 5.2 (Self-regulating positive gradient).
Define $\sigma^{+}\!=\sigma\!\bigl{(}\beta\operatorname{LR}^{+}\bigr{)}$. Then, we have:
$\partial_{\pi^{+}}\ell_{\mathrm{MIO}}=\frac{\beta}{\pi^{+}}\!\bigl{(}1.5\sigma^{+}-1\bigr{)}\;\begin{cases}>0&\text{if }\;\sigma^{+}>\tfrac{2}{3},\\[4.0pt] <0&\text{if }\;\sigma^{+}<\tfrac{2}{3}.\end{cases}$
Hence, MIO exhibits an intrinsic clamping mechanism: once a chosen token attains excessive probability ($\sigma^{+}\!>\!2/3$) the gradient turns positive, gently pushing its likelihood downward and mitigating over-fitting; conversely, if the token remains under-represented ($\sigma^{+}\!<\!2/3$) the gradient is negative, encouraging further amplification. This self-adjusting, bi-phase behaviour amounts to an implicit self-paced curriculum: optimisation effort is automatically reallocated from already-well-learned positives to the more informative, harder ones, yielding superior robustness and sample efficiency compared with DPO.
6 Conclusion
In this work, we propose a novel perspective that frames RLHF and DPO as a unified form of contrastive learning over the distributions of chosen and rejected responses, grounded in mutual information. Building on this connection, we theoretically show that RLHF does not extend the base model’s capacity but merely amplifies its existing capabilities. Furthermore, we introduce the Mutual Information Optimization (MIO), replacing the high-variance Donsker-Varadhan (DV) bound with a Jensen-Shannon (JS) bound. MIO demonstrates superior preservation of reasoning and mathematical abilities and offers greater training stability compared to existing baselines. Extensive experimental results demonstrate that MIO achieves superior performance on a wide range of mathematical and reasoning benchmarks. We hope that our work could offer useful insights for the community and help address the limitations of current RLHF methods. This work does not explore the effects of alternative mutual information estimators, which we leave as a promising direction for future research.
Appendix A Searching for a critic $T_{\phi}$ is equivalent to training a reward model.
Energy–based form of the chosen model.
Throughout the analysis we assume that the preference-filtered policy $\pi_{\text{chosen}}$ is obtained from a frozen reference model $\pi_{\text{ref}}$ via an energy re-weighting:
$\pi_{\text{chosen}}(y\mid x)\;=\;\pi_{\text{ref}}(y\mid x)\,\frac{e^{\alpha\,r(x,y)}}{Z(x)},$ (A.1)
where $r(x,y)$ is a reward signal, $\alpha>0$ a temperature, and $Z(x)=\mathbb{E}_{y\sim\pi_{\text{ref}}}e^{\alpha r(x,y)}$ the normaliser.
Definition A.1.
For a positive pair $(x,y_{w})$ we define
$T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})$ $=\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{chosen}}(y_{w}\mid x)},$ (A.2)
$r(x,y_{w})$ $=\beta\!\Bigl{[}\log\tfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}+\log Z(x)\Bigr{]},$ (A.3)
with a scaling constant $\beta>0$.
Substitute (A.3) into (A.1).
$\pi_{\text{chosen}}(y_{w}\mid x)=\pi_{\text{ref}}^{\,1-\alpha\beta}\;\pi_{\theta}^{\,\alpha\beta}\;Z(x)^{\,\alpha\beta-1}.$ (A.4)
Insert (A.4) into (A.2).
$T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})=(1-\alpha\beta)\Bigl{[}\log\pi_{\theta}(y_{w}\mid x)-\log\pi_{\text{ref}}(y_{w}\mid x)+\log Z(x)\Bigr{]}.$ (A.5)
Which is to say:
$T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})=\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{chosen}}(y_{w}\mid x)}\propto\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}$ (A.6)
Compact form.
Introducing a single scaling parameter
$\gamma\;\coloneqq\;\frac{1-\alpha\beta}{\beta},$ (A.7)
and using (A.3) in (A.5) we arrive at
$\boxed{\;T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})\;=\;\gamma\,r(x,y_{w})\;}$ (A.8)
for the positive pair, and analogously
$\boxed{\;T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})\;=\;\gamma\,r(x,y_{l})\;}$ (A.9)
for the negative pair $(x,y_{l})$.
Appendix B Maximizing mutual information towards the positive model inherently distances it from the negative model.
Definition B.1.
For every prompt $x$ we draw $y_{w}\sim\pi_{\text{chosen}}(\cdot\!\mid\!x)$ (positive) and $y_{l}\sim\pi_{\text{rejection}}(\cdot\!\mid\!x)$ (negative). Fix a critic $T_{\phi}\!\in\!\mathcal{F}$. With $z^{+}=T_{\phi}(x,y_{w}),\;z^{-}=T_{\phi}(x,y_{l})$, define the two DV / InfoNCE estimators:
$\hat{I}^{+}$ $\;=\;\log\!\frac{e^{z^{+}}}{e^{z^{+}}\!+e^{z^{-}}}\;=\;-\log\!\bigl{(}1+e^{-(z^{+}-z^{-})}\bigr{)},$ (B.1)
$\hat{I}^{-}$ $\;=\;\log\!\frac{e^{z^{-}}}{e^{z^{+}}\!+e^{z^{-}}}\;=\;-\log\!\bigl{(}1+e^{z^{+}-z^{-}}\bigr{)}.$ (B.2)
Write the logit gap $\Delta\coloneqq z^{+}-z^{-}$. Using $\sigma(u)=\tfrac{1}{1+e^{-u}}$, we have $\hat{I}^{+}=\log\sigma(\Delta),\;\hat{I}^{-}=\log\sigma(-\Delta)$.
Gradients share the $\Delta$ direction but opposite sign.
$\frac{\partial\hat{I}^{+}}{\partial\Delta}$ $=\sigma(-\Delta),\qquad\frac{\partial\hat{I}^{-}}{\partial\Delta}=-\sigma(\Delta).$ (B.3)
Because $\sigma(u)>0$ for all $u\in\mathbb{R}$,
$\mathrm{sign}\!\Bigl{(}\partial_{\Delta}\hat{I}^{+}\Bigr{)}\;=\;-\;\mathrm{sign}\!\Bigl{(}\partial_{\Delta}\hat{I}^{-}\Bigr{)}.$
Coupled parameter gradients.
Let the model parameters be $\theta$ and assume $z^{+},z^{-}$ are differentiable in $\theta$. By the chain rule
$\nabla_{\theta}\hat{I}^{+}\;=\;\frac{\partial\hat{I}^{+}}{\partial\Delta}\,\nabla_{\theta}\Delta,\quad\nabla_{\theta}\hat{I}^{-}\;=\;\frac{\partial\hat{I}^{-}}{\partial\Delta}\,\nabla_{\theta}\Delta.$
Combine (B.3)–(B):
$\nabla_{\theta}\hat{I}^{-}\;=\;-\,\frac{\sigma(\Delta)}{\sigma(-\Delta)}\;\nabla_{\theta}\hat{I}^{+}.$
The scalar factor $-\sigma(\Delta)/\sigma(-\Delta)\,<0$; hence
[…]
i.e., the two gradient vectors always point in opposite directions.
Consequence.
Any update $\theta\leftarrow\theta+\eta\nabla_{\theta}\hat{I}^{+}\;(\eta>0)$ that increases the positive MI estimator must decrease the negative MI estimator:
$\hat{I}^{+}(\theta_{\text{new}})>\hat{I}^{+}(\theta),\;\Longrightarrow\;\hat{I}^{-}(\theta_{\text{new}})<\hat{I}^{-}(\theta).$
Hence, maximizing the estimated mutual information between $\pi_{\theta}$ and $\pi_{\text{chosen}}$ simultaneously minimizes the estimated mutual information between $\pi_{\theta}$ and $\pi_{\text{rejection}}$. This coupling holds for every step of gradient ascent and therefore for the final optimum, completing the proof.
Appendix C Mutual-Information Estimation: Optimal Critic and Practical Surrogate
C.1 DV / MINE Sampling Scheme for RLHF
Positive (joint) draw.
For each prompt $x\sim D$ we sample a chosen response $y_{w}\sim\pi_{\text{c}}(\cdot\mid x)$, defining the joint measure $dP_{\pi_{\theta},\pi_{\text{c}}}=D(x)\,\pi_{\text{c}}(y\mid x)$.
Negative (product) draw.
Conditioned on the same $x$ we draw a response $y\sim\bar{\pi}(\cdot\mid x)$ from the mixed policy $\bar{\pi}=\tfrac{1}{2}\pi_{\text{c}}+\tfrac{1}{2}\pi_{\text{r}},$ yielding the product measure $dP_{\pi_{\theta}}\,dP_{\bar{\pi}}$. Since $D(x)$ cancels in any density ratio, $\pi_{\theta}$ appears only as a marginal over the prompt.
Modified DV bound.
Under these measures the Donsker–Varadhan (DV) / MINE lower bound reads
$I(\pi_{\theta},\pi_{\text{chosen}})$ $=\sup_{\phi}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}[T_{\phi}]-\log(\mathbb{E}_{P_{\pi_{\theta}}}{\mathbb{E}_{P_{\pi_{\text{chosen}}}}}[e^{T_{\phi}}])$ (C.1)
$\geq\sup_{\phi}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}[T_{\phi}]-\log(\mathbb{E}_{P_{\pi_{\theta}}}\{{\mathbb{E}_{P_{\pi_{\text{chosen}}}}}[e^{T_{\phi}}]+{\mathbb{E}_{P_{\pi_{\text{rejection}}}}}[e^{T_{\phi}}]\})$ (C.2)
$=\sup_{\phi}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}[T_{\phi}]-\log(\mathbb{E}_{P_{\pi_{\theta}}}\{2{\mathbb{E}_{P_{\bar{\pi}}}}[e^{T_{\phi}}]\})$ (C.3)
$=\sup_{\phi}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}[T_{\phi}]-\log(\mathbb{E}_{P_{\pi_{\theta}}P_{\bar{\pi}}}[e^{T_{\phi}}])-\log 2$ (C.4)
$=\sup_{\phi}\mathbb{E}_{P_{\pi_{\theta}\pi_{\text{chosen}}}}\log e^{[T_{\phi}]}-\log(\mathbb{E}_{P_{\pi_{\theta}}P_{\bar{\pi}}}[e^{T_{\phi}}])-\log 2$ (C.5)
C.2 Optimal Critic for the Mixed-product Bound
The optimum is a log-density ratio.
Equation (C.5) is a Fenchel–Legendre dual:
$\sup_{T}\bigl{[}\mathbb{E}_{P_{\text{joint}}}[T]-\log\mathbb{E}_{P_{\text{prod}}}[e^{T}]\bigr{]}\;=\;\mathrm{KL}\!\bigl{(}P_{\text{joint}}\,\|\,P_{\text{prod}}\bigr{)},$
whose supremum is attained when the Fenchel–Young equality holds, i.e.
$T^{\star}(x,y)\;=\;\log\frac{dP_{\text{joint}}}{dP_{\text{prod}}}(x,y).$
Substituting $P_{\text{joint}}\!=\!P_{\pi_{\theta},\pi_{\text{c}}}$ and $P_{\text{prod}}\!=\!P_{\pi_{\theta}}P_{\bar{\pi}}$ gives
$\boxed{T^{\star}(x,y)=\log\frac{\pi_{\text{c}}(y\mid x)}{\bar{\pi}(y\mid x)}}.$ (C.2)
Intuitively, $T^{\star}$ rewards samples that are more likely under $\pi_{\text{c}}$ than under the negative pool $\bar{\pi}$.
C.3 Why $T^{\star}$ Is Impractical
Evaluating $\bar{\pi}(y\mid x)$ requires two forward passes ($\pi_{\text{c}}$ and $\pi_{\text{r}}$) for every negative candidate, multiplying GPU cost. Moreover, $e^{T^{\star}}=\pi_{\text{c}}/\bar{\pi}$ can be extremely heavy-tailed when $\pi_{\text{r}}$ diverges from $\pi_{\text{c}}$, rendering the Monte-Carlo estimate of $\mathbb{E}[e^{T^{\star}}]$ numerically unstable. For these reasons optimising (C.2) is unattractive in practice.
C.4 A Restricted Critic Family
If we therefore constrain the critic to the log-ratio sub-family
$\boxed{\mathcal{F}_{\text{sub}}=\bigl{\{}T_{\theta,c}(x,y)=\log\tfrac{\pi_{\theta}(y\mid x)}{\pi_{\text{c}}(y\mid x)}+c\;\bigm{|}\;c\in\mathbb{R}\bigr{\}},}$ (C.3)
whose evaluation requires only one forward pass through $\pi_{\text{c}}$.
Plugging $T_{\theta,c}$ into (C.5) yields the surrogate bound
$\hat{I}(\theta):=\sup_{c\in\mathbb{R}}\Bigl{[}\mathbb{E}_{P_{\pi_{\theta},\pi_{\text{c}}}}[T_{\theta,c}]-\log\mathbb{E}_{P_{\pi_{\theta}}P_{\bar{\pi}}}[e^{T_{\theta,c}}]\Bigr{]}-\log 2,$ (C.4)
which satisfies
$\boxed{\hat{I}(\theta)\;\leq\;I^{\star}_{\bar{\pi}}(\theta)\;\leq\;I(\pi_{\theta},\pi_{\text{c}})}.$ (C.5)
The first inequality reflects the restriction $\mathcal{F}_{\text{sub}}\!\subset\!\mathcal{F}_{\text{all}}$; the second arises from replacing $\pi_{\text{c}}$ by $\bar{\pi}$ in the denominator. Despite this looseness, Proposition B.1 in the main text shows that the positive and negative gradients are always anti-parallel, guaranteeing a contrastive signal at every optimisation step.
Appendix D Upper Bounds on the Jensen Gap
We consider the approximation
$\mathbb{E}_{\mathrm{J}}\!\bigl{[}T_{\phi}\bigr{]}\;-\;\log\mathbb{E}_{\pi_{\theta}}\!\Bigl{[}\tfrac{\pi_{\theta}}{\pi_{\mathrm{chosen}}}\Bigr{]}\;\approx\;\mathbb{E}_{\mathrm{J}}\!\bigl{[}T_{\phi}\bigr{]}\;-\;\mathbb{E}_{\pi_{\theta}}\!\Bigl{[}\log\tfrac{\pi_{\theta}}{\pi_{\mathrm{chosen}}}\Bigr{]},$
and denote the resulting error (Jensen gap) by
$\Delta\;=\;\mathbb{E}_{\pi_{\theta}}\!\bigl{[}\log f(Y)\bigr{]}\;-\;\log\mathbb{E}_{\pi_{\theta}}\!\bigl{[}f(Y)\bigr{]},\qquad f(Y)=\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{chosen}}(y\mid x)}.$
The variance-based bound.
Let $\mu=\mathbb{E}_{\pi_{\theta}}[f]$ and $\sigma^{2}=\mathrm{Var}_{\pi_{\theta}}(f).$ If $\sigma/\mu\ll 1$, then by a second-order Taylor expansion of $\log$ around $\mu$,
$\Delta=\log\mu-\mathbb{E}\bigl{[}\log f\bigr{]}\;\leq\;\frac{\sigma^{2}}{2\,\mu^{2}}\,.$
As the fluctuations of $f$ decrease (i.e. $\sigma^{2}\!\to\!0$), the Jensen gap $\Delta$ approaches zero. Because $\pi_{\mathrm{chosen}}$ is an energy-based model,
$f(Y)=\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{chosen}}(y\mid x)}\;\propto\;\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}+\log Z(x).$
That is to say, in order to curb the additional Jensen-gap error we must keep the ratio $\tfrac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}$ close to unity—i.e., constrain the policy’s deviation from the reference distribution—thereby limiting $\mathrm{Var}(f)$ and tightening the bound; this perspective provides an alternative justification for the ratio-clipping heuristic employed by PPO-style methods.
Appendix E Supplementary: Preliminaries
Notation
Let $x\!\in\!\mathcal{X}$ be a user prompt, $y\!\in\!\mathcal{Y}$ a textual response, and $\pi(y\mid x)$ a conditional language model (policy). We denote
  • $\pi_{\text{ref}}$ –– a fixed reference model (e.g. SFT checkpoint),
  • $\pi_{\theta}$ –– the trainable policy with parameters $\theta$,
  • $\pi_{\text{chosen}}$ — the human-preferred generator producing positive samples $y_{w}$, modeled as an energy-based model (see Equation A.1).
  • $\pi_{\text{rejection}}$ — the lower-preference generator producing negative samples $y_{l}$, also instantiated as an energy-based model.
  • $\bar{\pi}=\tfrac{1}{2}(\pi_{\text{chosen}}+\pi_{\text{rejection}})$ –– the mixed negative pool used in MINE sampling.
Reinforcement Learning from Human Feedback.
Given the estimated reward function $r(\mathbf{x},\mathbf{y})$, dictating the human preferences, RLHF fine-tunes policy $\pi_{\theta}$ by optimizing the following objective:
$\max_{\pi_{\theta}}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\mathbf{y}|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]-\beta\mathrm{D}_{\mathrm{KL}}\left[\pi_{\theta}(\mathbf{y}|\mathbf{x})\|\pi_{\mathrm{ref}}(\mathbf{y}|\mathbf{x})\right],$ (C.6)
where $\beta>0$ is an appropriate KL penalty coefficient. RLHF typically optimizes the above objective in Equation C.6 using RL algorithms, such as PPO [49]. Although RLHF with PPO has achieved remarkable success, the training process of PPO is unstable because of the high variance of the estimates of the policy gradients [50].
Reward Modeling.
One standard approach to reward modeling is to fit a reward function $r_{\phi}(\mathbf{x},\mathbf{y})$ with the BT preference model in Equation (1). Specifically, the reward function $r_{\phi}(\mathbf{x},\mathbf{y})$ can be estimated by maximizing the log-likelihood over preference feedback $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$:
$\mathcal{L}_{\mathrm{RM}}(\phi;\mathcal{D})=\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim\mathcal{D}}\left[-\log\sigma\left(r_{\phi}(\mathbf{x},\mathbf{y}_{w})-r_{\phi}(\mathbf{x},\mathbf{y}_{l})\right)\right].$ (C.7)
Mutual Information and the DV Lower Bound
For two random variables $(Y,Z)$ with joint $P_{YZ}$ and marginals $P_{Y},P_{Z}$, the mutual information is $I(Y;Z)=\mathrm{KL}(P_{YZ}\,\|\,P_{Y}P_{Z})$. The Donsker–Varadhan (DV) variational form writes
$I(Y;Z)=\sup_{T\in\mathcal{F}}\Bigl{(}\mathbb{E}_{P_{YZ}}[T]-\log\mathbb{E}_{P_{Y}P_{Z}}[e^{T}]\Bigr{)},$ (C.8)
where $\mathcal{F}$ is any function class with finite log-moment [21, 51]. Belghazi et al. [22] propose parametrising $T_{\phi}$ by a neural network and maximising the RHS of (C.8); the resulting estimator is known as MINE.
Energy-based Models.
Energy-based models (EBMs) [24] define the distribution through an energy function. For $\mathbf{y}\in\mathbb{R}^{D}$, its probability density can be expressed as follows:
$p_{\theta}(\mathbf{y})=\exp(-E_{\theta}(\mathbf{y}))/Z_{\theta}(\mathbf{y}),$ (C.9)
where $E_{\theta}(\mathbf{y}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ is the energy function, mapping the data point $\mathbf{y}$ to a scalar, and $Z_{\theta}(\mathbf{y})=\sum_{\mathbf{y}}\exp(-E_{\theta}(\mathbf{y}))$ is the unknown normalization constant [52].
Appendix F Supplementary: Experimental Details
The source code used in our experiments is available at: https://anonymous.4open.science/r/MIO-63E6/
F.1 Training Details
Table 2: The hyperparameter search space for the baselines.
Method Method
DPO  $\beta\in[0.01,0.05,0.1]$ IPO  $\tau\in[0.01,0.1,0.5,1.0]$
NCA  $\beta\in[0.01,0.10,0.15]$ SLiC  $\lambda\in[0.1,0.5,1.0,10.0]$, $\delta\in[0.1,0.5,1.0,2.0]$
KTO  $\lambda_{l}=\lambda_{w}=1.0$ $\beta\in[0.01,0.05,0.1]$ SimPO  $\beta\in[2.0,2.5]$ $\gamma\in[0.3,0.5,1.0,1.2,1.4,1.6]$
DIL  $\beta\in[0.8,1.0,1.2]$ ORPO  $\beta\in[0.01,0.05,0.1]$
We follow the general training configurations established in SimPO. During the supervised fine-tuning (SFT) stage, we use a learning rate of $2\times 10^{-5}$. For both SFT and preference optimization, we set the batch size to 128 and the maximum sequence length to 2048. A cosine learning rate scheduler is applied with a 10% warm-up ratio over a single epoch. We use the Adam optimizer [53] in all experiments.
Method-specific hyperparameters are selected according to the search strategy described in SimPO. Each baseline method uses its own dedicated hyperparameter set, as detailed in Table 2. The learning rate for each method is tuned within the range $\{3\mathrm{e}{-7},5\mathrm{e}{-7},7\mathrm{e}{-7},1\mathrm{e}{-6},5\mathrm{e}{-6}\}$.
To mitigate length bias, we normalize the response likelihood by computing the average log-probability of all tokens in the response under the policy model, following SimPO and DIL.
All training experiments are conducted using six NVIDIA A6000 GPUs (46GB each), with a total batch size of 384. We build on the publicly available alignment-handbook codebase.
F.2 The Details of Datasets
UltraFeedbackBinarized [54, 55] is a dataset comprising 64k prompts, each paired with four completions generated by various open-source and proprietary language models. GPT-4 assigns scores to these completions based on criteria such as helpfulness, honesty, and other qualitative metrics. To form binary preference pairs, the completion with the highest average score is selected as the preferred response, while one of the remaining three completions is randomly chosen as the rejected response.
F.3 Evaluation Tasks: Math and Reasoning Benchmarks
This section introduces the benchmark used for model evaluation. We follow prior work to evaluate the model fine-tuned on the UltraFeedbackBinarized dataset. Subsequently, we assess its performance on a suite of eight math- and reasoning-oriented tasks, aiming to comprehensively evaluate its quantitative reasoning ability, multi-step problem-solving skills, and domain-specific competence.
HendrycksMath [56] consists of 12,500 high school mathematics problems spanning algebra, geometry, and number theory. Each problem is accompanied by a step-by-step solution, targeting rigorous symbolic reasoning.
MinervaMath [57] evaluates quantitative reasoning over STEM domains using university-level mathematics, science, and engineering problems. It emphasizes complex problem solving in structured formats.
MultiMedQA [58] is a multi-task benchmark for medical question answering, aggregating datasets from professional exams, clinical research, and consumer health queries. It measures both factual correctness and clinical alignment.
MathQA [59] contains over 37k multiple-choice math word problems paired with symbolic execution programs. The dataset encourages interpretable problem solving using operation-based formalisms.
GSM8K [60] includes 8,500 elementary school math word problems requiring multi-step reasoning. It is widely used to benchmark arithmetic and logical problem-solving capabilities of language models. For this task, we report results using the exact match metric under the flexible-extract setting, which allows for variations in answer formatting.
AQuA-RAT [61] focuses on algebraic word problems from the AQuA-RAT dataset, requiring rationales and supporting steps. It is part of the broader AGIEval benchmark designed to test cognitive ability in foundation models. For this task, we use normalized accuracy (acc_norm) as the evaluation metric to account for formatting variability in symbolic responses.
MATH-Hard is a curated subset from the Hugging Face Open LLM Leaderboard that emphasizes challenging math problems, including multi-hop symbolic reasoning and advanced problem types.
MuSR [62] targets multistep soft reasoning problems, typically presented in narrative forms such as logic puzzles and commonsense tasks. It evaluates language models’ ability to track state, causality, and inference chains over extended contexts.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Seeing in the Section 6
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Seeing in the Appendix B
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: This paper introduce a new algorithm, the hyperparameter will be released in supply materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Seeing in the supplementary material
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Seeing in the Section 4
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Seeing in the Section 4
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Seeing in the supplementary material.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Seeing in Section 1
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Seeing in the Section 6.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: This work does not involve the release of models or datasets with high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: See Section 1
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not introduce any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [No]
  74. Justification: This work does not involve study participants or activities requiring IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: LLMs were only used for writing and editing purposes, and not as part of the core methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.